# Compatibility of Kazhdan and Brauer Homomorphism

### v2

@@ 5. Compatibility of Brauer homomorphism and Kazhdan isomorphism
In this section, we prove Theorem 1.1. Let us first recall the notations and terminologies from the prceeding sections.
@@
Fix a positive integer $m$. Let $F$ and $F^{\prime}$ be the non-Archimedean local fields such that $F$ is $m$-close to $F^{\prime}$. Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$, where $p$ is the residue characteristic of both $F$ and $F^{\prime}$. According to Lemma 4.2, there exists a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of degree $l$ and $E^{\prime}$ is $em$-close to $E$, where $e\in\{1,l\}$.
@@
Let $F,F^{\prime},E$ and $E^{\prime}$ be as above. Then for any split connected reductive group G defined over $\mathbb{Z}$, we have
${\rm Kaz}_{m}^{F}\circ\overline{\rm Br}=\overline{\rm Br}^{\prime}\circ{\rm Kaz}_{em}^{E}.$
@@
The proof is divided into two parts. In the first part, we consider the extension $E/F$ to be unramified. The second part deals with totally ramified case. To start with, let us recall that an arbitrary $h\in\mathcal{H}(G_{E},K_{E})^{\sigma}$ can be written as
@@ 5.1.2. Totally Ramified case
and it is possible only if $\xi^{l}=\mu$. Therefore, it follows that
$\pi_{\xi^{l}}\in K_{E}c_{i}^{-1}a_{i}\pi_{\xi^{l}}a_{j}^{-1}c_{j}K_{E}$
@@
$(\overline{c_{i}}^{-1}\overline{a_{i}},\overline{a_{j}}^{-1}\overline{c_{j}})\in\Gamma_{\xi^{l}}(E).$
Under the bijection $\Gamma_{\xi^{l}}(E)\rightarrow\Gamma_{\xi^{l}}(E^{\prime})$, we have
$(\overline{c_{i}}^{-1}\overline{a_{i}},\overline{a_{j}}^{-1}\overline{c_{j}})\in\Gamma_{\xi^{l}}(E^{\prime})$
@@
$(\overline{c_{i}}^{\prime-1}\overline{a_{i}}^{\prime},\overline{c_{j}}^{\prime-1}\overline{a_{j}}^{\prime})=(\overline{c_{i}^{\prime}}^{-1}\overline{a_{i}^{\prime}},\overline{c_{j}^{\prime}}^{-1}\overline{a_{j}^{\prime}})\in\Gamma_{\xi^{l}}(E^{\prime}),$
where $(c_{i}^{\prime},c_{j}^{\prime})\in T_{\xi^{l}}(F^{\prime})$. Thus we have
@@ 6. Linkage over close local fields
In this section, an application of Theorem 5.1 is given in the context of linkage. To be more precise, we show that the linkage is compatible with Kazhdan isomorphism. To prove this, we first give a formulation of the compatibility of linkage with Brauer homomorphism, as described in [TV16, Section 6.2], and then we use Theorem 5.1 to get the required compatibility. To start with, we first recall the Tate cohomology and the concept of linkage. For a reference, see [TV16, Section 3 and 6].
@@ 6.1. Tate cohomology
Let $\Gamma$ be a cyclic group of order $n$ with generator $\gamma$. Let $M$ be an abelian group with an action of $\Gamma$. Let $T_{\gamma}$ be the automorphism of $M$ defined by
@@
Let $N={\rm id}+T_{\gamma}+T_{\gamma^{2}}+\dots+T_{\gamma^{n-1}}$ be the norm operator. The Tate cohomology groups $\widehat{H}^{0}(M)$ and $\widehat{H}^{1}(M)$ are defined as
@@
6.1.1. Tate Cohomology of sheaves on $\ell$-spaces
Let $X$ be an $\ell$-space (i.e., Hausdorff, locally compact and totally disconnected) with an action of a finite group $\langle\gamma\rangle$ of prime order $l$. Let $\mathcal{F}$ be a sheaf of $k$-modules on $X$. If $\mathcal{F}$ is $\gamma$-equivariant, then $\gamma$ can be regarded as a map of restricted sheaves $\mathcal{F}|_{X^{\gamma}}\rightarrow\mathcal{F}|_{X^{\gamma}}$ and the Tate cohomology is defined as
@@ 6.1.1. Tate Cohomology of sheaves on $\ell$-spaces
For each $i\in\{0,1\}$, the object $\widehat{H}^{i}(\mathcal{F}|_{X^{\gamma}})$ is a sheaf on $X^{\gamma}$.
@@ 6.2. Linkage
Let $F$ be a non-archimedean local field and let $G$ be the $F$-points of a connected reductive group defined over $F$. A representation $(\rho,V)$ of $G$ is called smooth if for every vector $v\in V$, the $G$-stabilizer of $v$ is an open subgroup of $G$. Moreover, such a representation $\rho$ is called admissible if for each compact open subgroup $K$ of $G$, the space of $K$-fixed vectors $V^{K}$ (sometimes denoted by $\rho^{K}$) is finite-dimensional over $k$. In this subsection, all the representations are assumed to be admissible and the representation spaces are defined over $k$. We denote by $\mathcal{R}(G)$ the category of smooth admissible $k$-representations of $G$.
@@ 6.2.1. Frobenius Twist
Then the action of $G$ on $V$ induces a representation $\rho^{(l)}$ of $G$ on $V^{(l)}$. The representation $(\rho^{(l)},V^{(l)})$ is called the Frobenius twist of the representation $(\rho,V)$.
@@
Let $\sigma$ be an automorphism of $G$ of prime order $l$. An irreducible representation $(\Xi,W)$ of $G$ is called $\sigma$-fixed if $\Xi$ is isomorphic to $\Xi\circ\sigma$. For such representation $\Xi$, there is an unique action of $\sigma$ on the space $W$, which is compatible with the action of $\sigma$ on $G$ ([TV16, Proposition 6.1]). Then we have the Tate cohomology space $\widehat{H}^{i}(W)$, defined as $k$-representation of $G^{\sigma}$. We denote this representation by $\widehat{H}^{i}(\Xi)$.
@@
An irreducible representation $\rho$ of $G^{\sigma}$ is linked with the representation $\Xi$ if the Frobenius-twist $\rho^{(l)}$ occurs as a Jordan-Holder constituent of $\widehat{H}^{0}(\Xi)$ or $\widehat{H}^{1}(\Xi)$.
6.2.3.
It was conjectured by Venkatesh-Treumann ([TV16, Section 6.3]) that $\widehat{H}^{i}(\Xi)$ is of finite length as a representation of $G^{\sigma}$. We give a proof of this conjecture for the group ${\rm GL}_{n}$, under the following set up.
[…]
We start with a non-archimedean local field $L$. Let $P_{n}(L)$ and $U_{n}(L)$ be the subgroups of ${\rm GL}_{n}(L)$, given by
@@
respectively. We denote by $X_{L}$ the coset space $P_{n}(L)/U_{n}(L)$. Let $\theta_{n}$ be the character of $U_{n}(L)$, defined as
$\theta_{n}((u)_{i,j}):=\psi_{L}(u_{n-1,n})$, for $(u)_{i,j}\in U_{n}(L)$.
We have four fundamental functors:
@@
given by $\Psi^{-}=r_{U_{n}(L),{\rm id}}$, $\Phi^{-}=r_{U_{n}(L),\theta_{n}}$, $\Psi^{+}=i_{U_{n}(L),{\rm id}}$, and $\Phi^{+}=i_{U_{n}(L),\theta_{n}}$.
@@
where $\tau_{m}=(\Phi^{+})^{m-1}(\Phi^{-})^{m-1}(\tau)$ and $\tau_{m}/\tau_{m+1}=(\Phi^{+})^{m-1}(\Psi^{+})(\tau^{(m)})$. Using these, we now deduce:
@@
Let $\Xi$ be an irreducible $\sigma$-fixed representation of ${\rm GL}_{n}(E)$. Let $T_{\sigma}:\Xi\rightarrow\Xi\circ\sigma$ be an isomorphism with $T_{\sigma}^{l}={\rm id}$. Then the Tate cohomology $\widehat{H}^{i}(\Xi)$ with respect to the operator $T_{\sigma}$ is a finite length representation of ${\rm GL}_{n}(F)$.
@@
We use induction on the integer $n$. Assume that the proposition is true for all positive integers less than $n$. Since $\psi_{E}(\sigma(x))=\psi_{E}(x)$, for $x\in E$, we get the isomorphism
@@
The map $T_{\sigma}$ together with (6.1) and (6.2) induces an isomorphism between $\Xi^{(m)}$ and $(\Xi^{(m)})\circ\sigma$, and also between the representations $\Xi_{m}$ and $\Xi_{m}\circ\sigma$ for $m\geq 1$. Using induction hypothesis, we get that the representation $\widehat{H}^{i}(\Xi^{(m)})$ is a finite length representation of ${\rm GL}_{n-m}(F)$. Since $P_{n}(E)$ is stable under the action of $\sigma$, for any $n\geq 1$, we get that there is a bijection between the sets $X_{E}^{\sigma}$ and $X_{F}$. Then according to [TV16, Proposition 3.3.1], we have
$\widehat{H}^{i}(\Xi_{m}/\Xi_{m+1})=(\Phi^{+})^{m-1}(\Psi^{+})\big{(}\widehat{H}^{i}(\Xi^{(m)})\big{)}.$
Hence, for all $m\geq 1$, the representation $\widehat{H}^{i}(\Xi_{m}/\Xi_{m+1})$ is finite length representation of $P_{n}(F)$. Now, using spectral sequence induced by canonical filtration of ${\rm Res}_{P_{n}(E)}(\Xi)$, we get that $\widehat{H}^{i}(\Xi)$ is finite length representation of $P_{n}(F)$, and this implies that $\widehat{H}^{i}(\Xi)$ is finite length representation of ${\rm GL}_{n}(F)$. ∎
It is reasonable to expect that Proposition 6.2 holds true for any general connected reductive algebraic group G over $F$. In the next section, we give a formulation of the compatibility of Brauer homomorphism with linkage, under the finite length assumption of the Tate cohomology $\widehat{H}^{i}(\Xi)$ for any irreducible $\sigma$-fixed representation $\Xi$ of $G=\textbf{G}(F)$. Then we show that this formulaton is compatible with Kazhdan isomorphism.

### v4

@@ 5. Compatibility of Brauer homomorphism and Kazhdan isomorphism
In this section, we prove Theorem 1.1. Let us first recall the notations and terminologies from the preceding sections.
@@
Fix a positive integer $m$. Let $F$ and $F^{\prime}$ be the non-Archimedean local fields such that $F$ is $m$-close to $F^{\prime}$. Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$, where $p$ is the residue characteristic of both $F$ and $F^{\prime}$. According to Lemma 4.6, there exists a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of degree $l$ and $E^{\prime}$ is $em$-close to $E$, where $e=1$ if $E/F$ is unramified and $e=l$ if $E/F$ is totally ramified.
@@
Let $F,F^{\prime},E$ and $E^{\prime}$ be as above. Then, for any connected split reductive group G defined over $\mathbb{Z}$, we have
${\rm Kaz}_{m}^{F}\circ\overline{\rm Br}=\overline{\rm Br}^{\prime}\circ{\rm Kaz}_{em}^{E},$
where $e=1$ if $E/F$ is unramified and $e=l$ if $E/F$ is totally ramified.
@@
We divide the proof into two parts. In the first part, we consider the extension $E/F$ unramified. The second part deals with the totally ramified case. To begin with, let us recall that an arbitrary $h\in\mathcal{H}(G_{E},K_{E})^{\sigma}$ can be written as
@@ 5.1.2. Totally Ramified case
and it is possible only if $l\xi=\mu$. Therefore, it follows that
$\pi_{l\xi}\in K_{E}c_{i}^{-1}a_{i}\pi_{l\xi}a_{j}^{-1}c_{j}K_{E}$
@@
$(\overline{c_{i}}^{-1}\overline{a_{i}},\overline{a_{j}}^{-1}\overline{c_{j}})\in\Gamma_{l\xi}(E).$
Under the bijection $\Gamma_{l\xi}(E)\rightarrow\Gamma_{l\xi}(E^{\prime})$, we have
$(\overline{c_{i}}^{-1}\overline{a_{i}},\overline{a_{j}}^{-1}\overline{c_{j}})\in\Gamma_{l\xi}(E^{\prime})$
@@
$(\overline{c_{i}}^{\prime-1}\overline{a_{i}}^{\prime},\overline{c_{j}}^{\prime-1}\overline{a_{j}}^{\prime})=(\overline{c_{i}^{\prime}}^{-1}\overline{a_{i}^{\prime}},\overline{c_{j}^{\prime}}^{-1}\overline{a_{j}^{\prime}})\in\Gamma_{l\xi}(E^{\prime}),$
where $(c_{i}^{\prime},c_{j}^{\prime})\in T_{l\xi}(F^{\prime})$. Thus we have
@@ 6. Linkage over close local fields
This section gives an application of Theorem 5.1 in the context of representation theory. In the article [TV16], Treumann–Venkatesh defined the notion of linkage, which is conjectured to be compatible with the Langlands functorial transfer. Using Theorem 5.1, we show that linkage is compatible under close local fields. To prove this, we first formulate the compatibility of linkage with Brauer homomorphism, as described in [TV16, Section 6.2]. To begin with, we recall the notion of linkage and Tate cohomology. For precise reference, see [TV16, Section 3 and 6].
@@ 6.1. Tate cohomology
Let $\Gamma$ be a cyclic group of order $n$ with a generator $\gamma$, and let $M$ be an abelian group with an action of $\Gamma$. Let $T_{\gamma}$ be the automorphism of $M$ defined by
@@
Let $N={\rm id}+T_{\gamma}+T_{\gamma^{2}}+\dots+T_{\gamma^{n-1}}$ be the norm operator. The Tate cohomology groups $\widehat{H}^{i}(M)$, for $i\in\{0,1\}$, are defined as
@@
6.1.1. Tate Cohomology of sheaves on $l$-spaces
Let $X$ be an $l$-space (i.e., locally compact, totally disconnected and Hausdorff) with an action of a finite group $\langle\gamma\rangle$ of prime order $l$. Let $\mathcal{F}$ be a sheaf of $k$-modules on $X$. If $\mathcal{F}$ is $\gamma$-equivariant, then $\gamma$ can be regarded as a map of restricted sheaves $\mathcal{F}|_{X^{\gamma}}\rightarrow\mathcal{F}|_{X^{\gamma}}$ and the Tate cohomology is defined as
@@ 6.1.1. Tate Cohomology of sheaves on $l$-spaces
For each $i$, the object $\widehat{H}^{i}(\mathcal{F}|_{X^{\gamma}})$ is a sheaf on $X^{\gamma}$. If $\Gamma_{c}(X;\mathcal{F})$ denotes the space of compactly supported sections on $\mathcal{F}$, then we have the following result.
Proposition 6.1.
[TV16, Proposition 3.3.1] For each $i\in\{0,1\}$, the restriction map induces an isomorphism
$\widehat{H}^{i}(\Gamma_{c}(X;\mathcal{F}))\xrightarrow{\sim}\Gamma_{c}(X^{\gamma};\widehat{H}^{i}(\mathcal{F}))$.
@@ 6.2. Linkage
Let $F$ be a non-Archimedean local field, and $G$ be the $F$-points of a connected reductive group defined over $F$. A representation $(\rho,V)$ of $G$ is called smooth if for every vector $v\in V$, the $G$-stabilizer of $v$ is an open subgroup of $G$. The representation $\rho$ is called admissible if, for each compact open subgroup $K$ of $G$, the space of $K$-fixed vectors $V^{K}$ (sometimes denoted by $\rho^{K}$) is finite-dimensional over $k$. In this subsection, all the representations are assumed to be admissible, and the representation spaces are defined over $k$. We denote by $\mathcal{R}(G)$ the category of smooth, admissible, $k$-representations of $G$.
@@ 6.2.1. Frobenius Twist
Then, the action of $G$ on $V$ induces a smooth representation $\rho^{(l)}$ of $G$ on the space $V^{(l)}$. The representation $(\rho^{(l)},V^{(l)})$ is called the Frobenius twist of $(\rho,V)$.
@@
Let $\sigma$ be an automorphism of $G$ of prime order $l$. An irreducible representation $(\Xi,W)$ of $G$ is called $\sigma$-fixed if $\Xi$ is isomorphic to the twisted representation $\Xi\circ\sigma$. In such a case, there is a unique action of $\sigma$ on $W$, compatible with the action of $\sigma$ on $G$ (see [TV16, Proposition 6.1]). Then the Tate cohomology groups $\widehat{H}^{i}(W)$, for $i\in\{0,1\}$, are defined as $k$-representations of $G^{\sigma}$. We denote these representations by $\widehat{H}^{i}(\Xi)$.
@@
An irreducible representation $\rho$ of $G^{\sigma}$ is linked with the representation $\Xi$ if the Frobenius twist $\rho^{(l)}$ occurs as a Jordan-Holder constituent of $\widehat{H}^{0}(\Xi)$ or $\widehat{H}^{1}(\Xi)$.
In [TV16, Section 6.3], the authors made a conjecture that the Tate cohomology groups $\widehat{H}^{i}(\Xi)$, $i\in\{0,1\}$, are of finite length as representations of $G^{\sigma}$. We give a proof of this conjecture for the general linear group ${\rm GL}_{n}$ in the local base change setting, using Bernstein–Zelevinksy derivatives.
6.3. Finiteness of Tate cohomology
Before going into the main theorem, we first recall the notion of derivatives for smooth representations of ${\rm GL}_{n}(L)$, where $L$ is a non-Archimedean local field. We refer to [Vig96, Chapter III] and [BZ77, Section 3] for more details. Let $P_{n}(L)$ and $U_{n}(L)$ be the subgroups of ${\rm GL}_{n}(L)$, given by
@@ 6.3. Finiteness of Tate cohomology
respectively. We use the short notation $X_{L,n}$ to denote the coset space $P_{n}(L)/P_{n-1}(L)U_{n}(L)$. Let $\psi_{L}:L\rightarrow k^{\times}$ be a non-trivial additive character. Let $N_{n}(L)$ be the group of all unipotent upper triangular matrices in ${\rm GL}_{n}(L)$. Let $\Theta_{L}$ be the character of $N_{n}(L)$, defined as
$\Theta_{L}((a_{ij})_{i,j=1}^{n}):=\psi_{L}(a_{12}+a_{23}+\cdots a_{n-1,n}),$ (6.1)
for $(a_{ij})_{i,j=1}^{n}\in N_{n}(L)$. By abuse of notation, the restriction ${\rm res}_{U_{n}(L)}\Theta_{L}$ is also denoted by $\Theta_{L}$. We have four fundamental functors:
@@
given by $\Psi^{-}=r_{U_{n}(L),{\rm id}}$, $\Phi^{-}=r_{U_{n}(L),\Theta_{L}}$, $\Psi^{+}=i_{U_{n}(L),{\rm id}}$, and $\Phi^{+}=i_{U_{n}(L),\Theta_{L}}$.
6.3.1.
@@
where $\tau_{m}=(\Phi^{+})^{m-1}(\Phi^{-})^{m-1}(\tau)$ and $\tau_{m}/\tau_{m+1}=(\Phi^{+})^{m-1}(\Psi^{+})(\tau^{(m)})$. We now deduce:
Lemma 6.3.
Let $\rho$ be a finite length representation of $G_{t}(L)$, where $1\leq t<n$. Then, the $P_{n}(L)$-representation $(\Phi^{+})^{n-t-1}(\Psi^{+})(\rho)$ is also of finite length.
Proof.
This is an immediate consequence of [Vig96, Chapter III, Subsection 1.5] and the exactness of the functor $(\Phi^{+})^{n-t-1}(\Psi^{+})$. ∎
6.3.2.
Let $F$ be a non-Archimedean local field with residue characteristic $p$. Let $E/F$ be a finite Galois extension of prime degree $l$ with $l\neq p$. Let $\sigma$ be a generator of ${\rm Gal}(E/F)$. Fix a non-trivial additive character $\psi_{F}:F\rightarrow k^{\times}$. Let $\psi_{E}$ be the character $\psi_{F}\circ{\rm Tr}_{E/F}$, where ${\rm Tr}_{E/F}$ is the trace function. We have the non-degenerate characters $\Theta_{F}$ and $\Theta_{E}$ of $N_{n}(F)$ and $N_{n}(E)$ respectively, defined by (6.1). We now prove the following finiteness result of Tate cohomology:
@@
Let $\Xi$ be an irreducible $\sigma$-fixed representation of ${\rm GL}_{n}(E)$. Let $T_{\sigma}:\Xi\rightarrow\Xi\circ\sigma$ be an isomorphism with $T_{\sigma}^{l}={\rm id}$. Then the Tate cohomology group $\widehat{H}^{i}(\Xi)$, with respect to the operator $T_{\sigma}$, has finite length as a representation of ${\rm GL}_{n}(F)$.
@@
We prove the proposition using induction on $n$. The case $n=1$ is clear. So, we assume that the proposition is true for all irreducible $\sigma$-fixed representations of ${\rm GL}_{t}(E)$ and for all $t<n$. Now, we consider $\Xi$ as a representation of $P_{n}(E)$. Since $\psi_{E}(\sigma(x))=\psi_{E}(x)$, for $x\in E$, we get the isomorphism
@@
Using (6.2) and (6.3), and the isomorphism $T_{\sigma}$, we get an isomorphism between the representations $\Xi_{m}$ and $\Xi_{m}\circ\sigma$, and also between the representations $\Xi^{(m)}$ and $\Xi^{(m)}\circ\sigma$, for all $m\leq n$.
Recall that $X_{E,m}$ denotes the coset space $P_{m}(E)/P_{m-1}(E)U_{m}(E)$. Since $P_{m-1}(E)U_{m}(E)$ is a $\sigma$-stable subgroup of $P_{m}(E)$, we have the following long exact sequence of non-abelian cohomology ([Ser, Appendix, Proposition 1]):
$0\longrightarrow P_{m-1}(F)U_{m}(F)\longrightarrow P_{m}(F)\longrightarrow X_{E,m}^{\sigma}\longrightarrow H^{1}(\sigma,P_{m-1}(E)U_{m}(E))\longrightarrow H^{1}(\sigma,P_{m}(E))$ (6.4)
Consider the short exact sequence of non-abelian ${\rm Gal}(E/F)$-modules
$0\longrightarrow U_{m}(E)\longrightarrow P_{m-1}(E)U_{m}(E)\longrightarrow P_{m-1}(E)\longrightarrow 0$ (6.5)
Using Hilbert’s theorem 90, we get that the pointed sets $H^{1}(\sigma,U_{m}(E))$ and $H^{1}(\sigma,P_{m-1}(E))$ are trivial. Then, it follows from the long exact sequence of non-abelian cohomology corresponding to (6.5) that $H^{1}(\sigma,P_{m-1}(E)U_{m}(E))=0$. Hence, the long exact sequence (6.4) gives the equality $X_{E,m}^{\sigma}=X_{F,m}$. Now, using Proposition 6.1 repeatedly $(m-1)$-times, we get the $P_{n}(F)$-equivariant isomorphism
$\widehat{H}^{i}(\Xi_{m}/\Xi_{m+1})\simeq(\Phi^{+})^{m-1}(\Psi^{+})\big{(}\widehat{H}^{i}(\Xi^{(m)})\big{)}.$
By induction hypothesis, the ${\rm GL}_{n-m}(F)$-representation $\widehat{H}^{i}(\Xi^{(m)})$ is of finite length, for all $m<n$. In view of Lemma 6.3, it then follows from the above isomorphism that the $P_{n}(F)$-representation $\widehat{H}^{i}(\Xi_{m}/\Xi_{m+1})$ is of finite length, for all $m<n$.
Now, for each $m\in\{1,2,\dots,n-1\}$, consider the short exact sequence of $P_{n}(E)$-representations
$0\longrightarrow\Xi_{m+1}\longrightarrow\Xi_{m}\longrightarrow\Xi_{m}/\Xi_{m+1}\longrightarrow 0$
Since ${\rm Gal}(E/F)$ is cyclic, the corresponding long exact sequence of Tate cohomology gives the following diagram:
${\widehat{H}^{0}(\Xi_{m+1})}$${\widehat{H}^{0}(\Xi_{m})}$${\widehat{H}^{1}(\Xi_{m}/\Xi_{m+1})}$${\widehat{H}^{0}(\Xi_{m}/\Xi_{m+1})}$${\widehat{H}^{1}(\Xi_{m})}$${\widehat{H}^{1}(\Xi_{m+1})}$
We denote the above exact sequence by $S(m)$. Recall that $\Xi_{n}$ is the compactly induced representation ${\rm ind}_{N_{n}(E)}^{P_{n}(E)}(\Theta_{E})$, and using [TV16, Proposition 3.3.1], we get that $\widehat{H}^{0}(\Xi_{n})$ is isomorphic to the irreducible representation ${\rm ind}_{N_{n}(F)}^{P_{n}(F)}(\Theta_{F}^{l})$ and $\widehat{H}^{1}(\Xi_{n})=0$. In particular, the representation $\widehat{H}^{i}(\Xi_{n})$, for $i\in\{0,1\}$, is of finite length. Now, using the exact sequence $S(n-1)$ and the finiteness of $\widehat{H}^{i}(\Xi_{n-1}/\Xi_{n})$, we get that $\widehat{H}^{i}(\Xi_{n-1})$ is a finite length representation of $P_{n}(F)$. Again, using the finiteness of both the representations $\widehat{H}^{i}(\Xi_{n-1})$ and $\widehat{H}^{i}(\Xi_{n-2}/\Xi_{n-1})$, it follows from the exact sequence $S(n-2)$ that the representation $\widehat{H}^{i}(\Xi_{n-2})$ is of finite length. Thus, after a finite number of similar inductive steps, we finally get that $\widehat{H}^{i}(\Xi)$ is of finite length as a representation of $P_{n}(F)$, and hence of ${\rm GL}_{n}(F)$. This completes the proof. ∎
It is reasonable to expect that Proposition 6.4 holds for arbitrary connected reductive algebraic groups G over $F$. Next, we give a formulation of the compatibility of Brauer homomorphism with linkage under the finite length assumption of the Tate cohomology $\widehat{H}^{i}(\Xi)$ for any irreducible $\sigma$-fixed representation $\Xi$ of $G=\textbf{G}(F)$. It enables us to study linkage from the module theoretic point of view via Brauer homomorphism. Then, using Theorem 5.1, we show that linkage is compatible with the Kazhdan isomorphism.